# Orbital orders under magnetic fields in cubic PrIr2Zn20

### v2

@@ Abstract
We study orbital orders in quadrupolar heavy-fermion compound PrIr2Zn20 under magnetic fields on the basis of the Landau theory. Assuming Eg orbital (electric quadrupolar) orders in the cubic symmetry with the ordering wavenumber at the L points in the face-centered cubic lattice Brillouin zone as observed experimentally, we construct the Landau free energy and analyze it. We find that the unidentified high-temperature ordered phase under the magnetic field $\bm{H}\parallel[001]$ reported earlier can be interpreted as the consequence of the internal rotation in the orbital moments of f electrons at the Pr site. We discuss the phase diagram for various magnetic-field directions and also possible double-$\bm{q}$ quadrupolar orders in this system.
@@ II Model
We start to introduce the exchange interactions between the Eg orbital moments in Sec. II.1 and the Landau free energy for the order parameters at the L points in Sec. II.2. Mode coupling terms related to the order parameter at the L points and other degrees of freedom are discussed in Sec. II.5
@@ II.1 Exchange interaction
Let us denote Eg quadrupole moments at the position $\bm{r}=(x,y,z)$ in the diamond structure as $\bm{\phi}_{\bm{r}}^{\alpha}=(u_{\bm{r}}^{\alpha},v_{\bm{r}}^{\alpha})$, where $\alpha=$A or B is the sublattice index. We take one of the A sublattice site at the origin $\bm{r}=(0,0,0)$ and its pair B sublattice site in the unit cell at $\bm{r}+(1,1,1)/4$. $\bm{\phi}_{\bm{r}}^{\alpha}$ form a two-dimensional vector field as a classical variable. Here, $(u,v)$ represents $(u,v)\sim\{2z^{2}-x^{2}-y^{2},\sqrt{3}(x^{2}-y^{2})\}$, where $\bm{r}=(x,y,z)$ is the real space coordinate. The Fourier transform in the wavenumber ${\bm{k}}$ space is defined as
@@
Here, the summation for ${\bm{r}}$ runs over the A-sublattice sites in the diamond structure and $N$ is the number of the unit cell.
The exchange Hamiltonian in the diamond structure is in general written as
@@
where $H^{\alpha\beta}_{{\bm{k}}}$ is a $2\times 2$ matrix. We focus on the orders with the ordering wavenumber at the L points: ${\bm{k}}_{1}\equiv(\pi,\pi,\pi)$, ${\bm{k}}_{2}\equiv(-\pi,\pi,\pi)$, ${\bm{k}}_{3}\equiv(-\pi,-\pi,\pi)$, and ${\bm{k}}_{4}\equiv(\pi,-\pi,\pi)$. Since the exchange interactions at the L points are isotropic due to the symmetry, we retain such exchange interactions while neglect anisotropic ones. Although the latter exchange interactions are generally present and can influence the ordering vectors and the excitation energy of quadrupole wave, it is no effect on the transition temperature for the order at the L points. Thus, a model without anisotropic interactions can be a good approximation as a minimal model. In this section, we consider exchange interactions up to the fourth neighbor interactions, while we will simplify the model in the later sections. Hereafter, the lattice constant $a$ is set to $a=1$.
The next nearest-neighbor and the fourth-neighbor bonds are those between the A-A (B-B) sublattices. They are
$H^{AA}_{{\bm{k}}}$ $=[J_{2}\gamma_{2}({\bm{k}})+J_{4}\gamma_{4}({\bm{k}})]\hat{\sigma}_{0}$ (3)
Here, $\hat{\sigma}_{0}={\rm diag}(1,1)$ and $\hat{\sigma}^{x,z}$ is the Pauli matrix and we have followed the phase convention in Ref. [Hattori2014]. $J_{i}$’s are isotropic coupling constants. The expressions for $\gamma$’s are listed in Table 1.
The nearest-neighbor and the third-neighbor bonds are those between the A-B sublattices. For the nearest-neighbor interaction, there is no anisotropic term owing to the symmetry. The form is thus given as
$H^{AB}_{{\bm{k}}}$ $=\frac{1}{2}[J_{1}\gamma_{1}({\bm{k}})+J_{3}\gamma_{3}({\bm{k}})]\hat{\sigma}_{0}.$ (4)
We note that $\gamma_{1}$ and $\gamma_{3}$ are complex. From the neutron experimental data, the ordering patterns are consistent with those expected for $J_{1}>0$.
@@
Table 1: exchange interaction coefficients. The values of $\gamma({\bm{k}})$’s at ${\bm{k}}=(0,0,0),(0,0,2\pi)$, and $(\pi,\pi,\pi)$ are also shown. $\eta_{{\bm{k}}}\equiv e^{i(k_{x}+k_{y}+k_{z})/4}$, $\zeta^{\pm}_{{\bm{k}}}\equiv\pm 2(\cos\frac{k_{x}}{2}+\cos\frac{k_{y}}{2}+\cos\frac{k_{z}}{2})-3$. The mean-field second-order transition temperatures $T_{c}$’s are $T_{c}=2|J_{1}-3J_{3}|+6J_{4}$ for the L points, $T_{c}=4|J_{1}+3J_{3}|-12J_{2}-6J_{4}$ for the $\Gamma$ points, and $T_{c}=4J_{2}-6J_{4}$ for the X points.
notation expression $\Gamma(0,0,0)$ X(0,0,$2\pi$) L($\pi,\pi,\pi$)
$\gamma_{2}({\bm{k}})$ $2\left(\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}+\cos\frac{k_{y}}{2}\cos\frac{k_{z}}{2}+\cos\frac{k_{z}}{2}\cos\frac{k_{x}}{2}\right)$ $6$ $-2$ 00
$\gamma_{4}({\bm{k}})$ $\cos k_{x}+\cos k_{y}+\cos k_{z}$ $3$ $3$ $-3$
$\gamma_{1}({\bm{k}})$ $4\eta_{{\bm{k}}}\left(\cos\frac{k_{x}}{4}\cos\frac{k_{y}}{4}\cos\frac{k_{z}}{4}+i\sin\frac{k_{x}}{4}\sin\frac{k_{y}}{4}\sin\frac{k_{z}}{4}\right)$ $4$ 00 $-2$
$\gamma_{3}({\bm{k}})$ $4\eta_{{\bm{k}}}\left(\cos\frac{k_{x}}{4}\cos\frac{k_{y}}{4}\cos\frac{k_{z}}{4}\zeta^{+}_{{\bm{k}}}+i\sin\frac{k_{x}}{4}\sin\frac{k_{y}}{4}\sin\frac{k_{z}}{4}\zeta^{-}_{{\bm{k}}}\right)$ $12$ 00 $6$
@@ II.3 L point free energy
When squeeze out the L point components ${\boldsymbol{\phi}}^{\alpha}_{n}=(u_{n}^{\alpha},v_{n}^{\alpha})\equiv{\boldsymbol{\phi}}^{\alpha}_{{\bm{k}}_{n}}(n=1,2,3,4)$ with the short handed notation, one finds
$F_{\rm L}$ $=\frac{a_{\rm L}}{2}\sum_{\alpha,n}|{\boldsymbol{\phi}}_{n}^{\alpha}|^{2}+J_{\rm L}\sum_{n}{\boldsymbol{\phi}}^{A}_{n}\cdot{\boldsymbol{\phi}}^{B}_{n}+\frac{c}{4}\sum_{\alpha,n}|{\boldsymbol{\phi}}_{n}^{\alpha}|^{4}$
@@
$+\frac{c^{\prime\prime\prime}}{4}\sum_{\alpha}\sum_{m<n}\sum_{m^{\prime}<n^{\prime}}({\boldsymbol{\phi}}^{\alpha}_{m}\cdot{\boldsymbol{\phi}}^{\alpha}_{n})({\boldsymbol{\phi}}^{\alpha}_{m^{\prime}}\cdot{\boldsymbol{\phi}}^{\alpha}_{n^{\prime}}).$ (6)
Note that $b$ terms are absent owing to the wavnumber conservation and $(m,n,m^{\prime},n^{\prime})=(1,2,3,4)$, $(1,3,2,4)$, and $(1,4,2,3)$ in the last term in the right-hand side. The coupling constants are denoted as
@@
$F_{L}^{1(4)}=\frac{c}{4}\sum_{\alpha}|{\boldsymbol{\phi}}^{\alpha}|^{4}.$ (9)
@@ II.4 $\Gamma$ point free energy
Under finite magnetic fields, the uniform quadrupole moment ${\boldsymbol{\phi}}_{0}^{\alpha}\equiv{\boldsymbol{\phi}}_{{\bm{k}}=\bf 0}^{\alpha}$ is induced via the effective field $\tilde{\bm{h}}$ in the second order perturbation of the magnetic field $\bm{H}$. The free energy up to the fourth-order terms for the $\Gamma$ point modes are
$F_{\Gamma}$ $=\sum_{\alpha}\Big{(}-\tilde{\bm{h}}\cdot{\boldsymbol{\phi}}_{0}^{\alpha}+\frac{a_{\Gamma 0}}{2}|{\boldsymbol{\phi}}_{0}^{\alpha}|^{2}\Big{)}+J_{\Gamma 0}{\boldsymbol{\phi}}_{0}^{\rm A}\cdot{\boldsymbol{\phi}}_{0}^{\rm B},$
@@
$a_{\Gamma 0}$ $=T+12J_{2}+6J_{4},\quad J_{\Gamma 0}=4J_{1}+12J_{3},$ (11)
and the effective field is given by
@@
and $g=(g_{J}\mu_{\rm B})^{2}[7/(3E_{4})-1/E_{5}]$ given by the excited $\Gamma_{4,5}$ CEF energy $E_{4,5}$.Hattori and Tsunetsugu (2014) $\mu_{\rm B}$ is the Bohr magneton and $g_{J}$ is the Lande’s $g$ factor for the multiplet with the angular momentum $J$; $g_{J}=4/5$ for the PrIr2Zn20 with $J=4$.
@@ II.5 Mode coupling terms
We now discuss mode coupling terms consisting of the fields at the L points and those at $\Gamma$ points. The lowest order contributions come from the cubic terms as
@@
$=-3\lambda_{b}\sum_{\alpha,n}\phi_{0}^{\alpha}\big{(}\phi_{n}^{\alpha}\big{)}^{2}\cos(\theta^{\alpha}_{0}+2\theta^{\alpha}_{n}),$ (13)
Here we have parameterized
@@
and the coupling constant $\lambda_{b}$ as an independent parameter. If $\lambda_{b}$ is estimated by Eq. (5), $\lambda_{b}=b$. The next leading order terms are the fourth order and given by
@@
Again, the coupling constant $\lambda_{c}^{(1,2)}$ has been introduced as an independent parameter. If $\lambda_{c}^{(1,2)}$ is estimated by Eq. (5), $\lambda^{(1)}_{c}=\lambda_{c}^{(2)}=c$.
In principle, ${\boldsymbol{\phi}}_{0}^{\alpha}$ can couple with ${\boldsymbol{\phi}}_{{\bm{k}}}^{\alpha}$ in the form of $\sim{\boldsymbol{\phi}}_{0}^{\alpha}{\boldsymbol{\phi}}_{\bm{k}}^{\alpha}{\boldsymbol{\phi}}_{-{\bm{k}}}^{\alpha}$, ${\boldsymbol{\phi}}_{0}^{\alpha}{\boldsymbol{\phi}}_{0}^{\alpha}{\boldsymbol{\phi}}_{\bm{k}}^{\alpha}{\boldsymbol{\phi}}_{-{\bm{k}}}^{\alpha}$. In this paper, we retain the order parameter ${\boldsymbol{\phi}}_{n}^{\alpha}$ and the induced moment ${\boldsymbol{\phi}}_{0}^{\alpha}$ under magnetic fields since the other fields are not important for the phase transition. Some renormalization effects arising from the field at the X points will be discussed in Sec. IV and Appendix B.
@@ III Results: single-q orders
In this section, we carry out minimization of the free energy $F_{L}+F_{\Gamma}+\delta F_{3}+\delta F_{4}$ [Eqs. (6), (10), (13), and (13)] with the single-$\bm{q}$ assumptions. The purpose here is to understand how mode coupling effects affect the order parameter configuration. The results in this section will become bases for understanding the full phase diagram shown in Sec. IV where we sill take into account multiple-$\bm{q}$ configurations. Since we are interested in the phase transition owing to order parameter at the L point, we assume the second-order coefficients of the free energy at the other wavenumber are always positive. We first discuss the candidates of the stable configurations by analyzing the third and fourth order terms in $F_{\rm L}$. Then, we will show the numerical results of full temperature-field phase diagram.

### v3

@@ Abstract
We study orbital orders in quadrupolar heavy-fermion compound PrIr2Zn20 under magnetic fields on the basis of the Landau theory. Assuming $E_{g}$ orbital (electric quadrupolar) orders in the cubic symmetry with the ordering wavenumber at the L points in the face-centered cubic lattice Brillouin zone as observed experimentally, we construct the Landau free energy and analyze it. We find that the unidentified high-temperature ordered phase under the magnetic field $\bm{H}\parallel[001]$ reported earlier can be interpreted as a consequence of the internal rotation in the orbital moments of f electrons at the Pr site. We also discuss the phase diagram for other magnetic-field directions and possible double-$\bm{q}$ quadrupolar orders in this system.
@@ II Model
We start to introduce the exchange interactions between the $E_{g}$ orbital moments in Sec. II.1 and the Landau free energy for the order parameters at the L points in Sec. II.2. Mode coupling terms related to the order parameter at the L points and other degrees of freedom are discussed in Sec. II.5
@@ II.1 Exchange interaction
Let us denote $E_{g}$ quadrupole moments at the position $\bm{r}=(x,y,z)$ in the diamond structure as $\bm{\phi}_{\bm{r}}^{\alpha}=(u_{\bm{r}}^{\alpha},v_{\bm{r}}^{\alpha})$, where $\alpha=A$ or $B$ is the sublattice index. We take one of the $A$ sublattice site at the origin $\bm{r}=(x,y,z)=(0,0,0)$ and its pair $B$ sublattice site in the unit cell at $(\tfrac{1}{4},\tfrac{1}{4},\tfrac{1}{4})$, where we set the lattice constant $a=1$. $\bm{\phi}_{\bm{r}}^{\alpha}$ form a two-dimensional vector field as a classical variable. The two components are $u$ $\sim$ $2z^{2}$--$x^{2}$--$y^{2}$ and $v$ $\sim$ $\sqrt{3}$($x^{2}$--$y^{2}$). The Fourier transform in the wavenumber ${\bm{k}}$ space is defined as
@@
Here, the summation for ${\bm{r}}$ runs over the $A$-sublattice sites and $N$ is the number of the unit cell.
The exchange Hamiltonian is in general written as
@@
where $H^{\alpha\beta}_{{\bm{k}}}$ is a $2\times 2$ matrix. We focus on the orders with the ordering wavenumber at the L points: ${\bm{k}}_{1}\equiv(\pi,\pi,\pi)$, ${\bm{k}}_{2}\equiv(-\pi,\pi,\pi)$, ${\bm{k}}_{3}\equiv(-\pi,-\pi,\pi)$, and ${\bm{k}}_{4}\equiv(\pi,-\pi,\pi)$. Since the exchange interactions at the L points are isotropic due to the symmetry, we retain such exchange interactions while neglect anisotropic ones. Although the anisotropic ones are generally present and can influence the ordering vectors and the excitation energy of quadrupole wave, it has no effect on the transition temperature for the order at the L points. Thus, a model without anisotropic interactions can be a good starting model as a minimal one. In this section, we consider exchange interactions up to the fourth neighbor interactions, while we will simplify the model in the later sections.
The $2\times 2$ blocks in Eq. (2) are given by
$H^{AA}_{{\bm{k}}}$ $=[J_{2}\gamma_{2}({\bm{k}})+J_{4}\gamma_{4}({\bm{k}})]\hat{1},$ (3)
$H^{AB}_{{\bm{k}}}$ $=\frac{1}{2}[J_{1}\gamma_{1}({\bm{k}})+J_{3}\gamma_{3}({\bm{k}})]\hat{1}.$ (4)
Here, $\hat{1}$ is a 2$\times$2 identity matrix. $J_{i}$’s are isotropic coupling constants. The expressions for $\gamma$’s are listed in Table 1. Note that $\gamma_{1}({\bm{k}})$ and $\gamma_{3}({\bm{k}})$ are complex. From the neutron experimental data, the ordering patterns are consistent with those expected for $J_{1}>0$ Iwasa et al. (2017).
@@
Table 1: exchange interaction coefficients. The values of $\gamma({\bm{k}})$’s at ${\bm{k}}=(0,0,0),(0,0,2\pi)$, ${\bm{k}}_{1}=(\pi,\pi,\pi)$, and the other L points are also shown. Here, we take the unit cell convention of $A$: (0,0,0) and $B$: $(\tfrac{1}{4},\tfrac{1}{4},\tfrac{1}{4})$ and thus, the ${\bm{k}}_{1}$ is apparently inequivalent to the remaining ${\bm{k}}_{2,3,4}$. $\eta_{{\bm{k}}}\equiv e^{i(k_{x}+k_{y}+k_{z})/4}$, $\zeta^{\pm}_{{\bm{k}}}\equiv\pm 2(\cos\frac{k_{x}}{2}+\cos\frac{k_{y}}{2}+\cos\frac{k_{z}}{2})-3$. The mean-field second-order transition temperatures $T_{c}$’s are $T_{c}=2|J_{1}-3J_{3}|+6J_{4}$ for the L points, $T_{c}=4|J_{1}+3J_{3}|-12J_{2}-6J_{4}$ for the $\Gamma$ points, and $T_{c}=4J_{2}-6J_{4}$ for the X points.
notation expression $\Gamma(0,0,0)$ X(0,0,$2\pi$) L($\pi,\pi,\pi$) L(others)
$\gamma_{2}({\bm{k}})$ $2\left(\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}+\cos\frac{k_{y}}{2}\cos\frac{k_{z}}{2}+\cos\frac{k_{z}}{2}\cos\frac{k_{x}}{2}\right)$ $6$ $-2$ 00 00
$\gamma_{4}({\bm{k}})$ $\cos k_{x}+\cos k_{y}+\cos k_{z}$ $3$ $3$ $-3$ $-3$
$\gamma_{1}({\bm{k}})$ $4\eta_{{\bm{k}}}\left(\cos\frac{k_{x}}{4}\cos\frac{k_{y}}{4}\cos\frac{k_{z}}{4}+i\sin\frac{k_{x}}{4}\sin\frac{k_{y}}{4}\sin\frac{k_{z}}{4}\right)$ $4$ 00 $-2$ $2$
$\gamma_{3}({\bm{k}})$ $4\eta_{{\bm{k}}}\left(\cos\frac{k_{x}}{4}\cos\frac{k_{y}}{4}\cos\frac{k_{z}}{4}\zeta^{+}_{{\bm{k}}}+i\sin\frac{k_{x}}{4}\sin\frac{k_{y}}{4}\sin\frac{k_{z}}{4}\zeta^{-}_{{\bm{k}}}\right)$ $12$ 00 $6$ $-6$
@@ II.3 L point free energy
We now concentrate on the L point components ${\boldsymbol{\phi}}^{\alpha}_{n}=(u_{n}^{\alpha},v_{n}^{\alpha})\equiv{\boldsymbol{\phi}}^{\alpha}_{{\bm{k}}_{n}}(n=1,2,3,4)$ with the short handed notation, we find the part of the free energy consisting of only ${\boldsymbol{\phi}}^{\alpha}_{{\bm{k}}_{n}}$’s is
$F_{\rm L}$ $=\frac{a_{\rm L}}{2}\sum_{\alpha,n}|{\boldsymbol{\phi}}_{n}^{\alpha}|^{2}+J_{\rm L}\sum_{n}(-1)^{p_{n}}{\boldsymbol{\phi}}^{A}_{n}\cdot{\boldsymbol{\phi}}^{B}_{n}+\frac{c}{4}\sum_{\alpha,n}|{\boldsymbol{\phi}}_{n}^{\alpha}|^{4}$
@@
$+\frac{c^{\prime\prime\prime}}{4}\sum_{\alpha}\sum_{(m,n,m^{\prime},n^{\prime})}({\boldsymbol{\phi}}^{\alpha}_{m}\cdot{\boldsymbol{\phi}}^{\alpha}_{n})({\boldsymbol{\phi}}^{\alpha}_{m^{\prime}}\cdot{\boldsymbol{\phi}}^{\alpha}_{n^{\prime}}).$ (6)
Note that $b$ terms are absent owing to the wavnumber conservation and $(m,n,m^{\prime},n^{\prime})=(1,2,3,4)$, $(1,3,2,4)$, and $(1,4,2,3)$ in the last term in the right-hand side. The factor $p_{n}(n=1,2,3,4)$ represents the sign of exchange interaction between the $A$ and $B$ sublattices for four kinds of the L points. Since our unit cell consists of the bond parallel to ${\bm{k}}_{1}$, $p_{n}=0$ is distinct from the others $p_{2,3,4}=1$. The coupling constants are denoted as
@@
$F_{L}^{1\bm{q}(4)}=\frac{c}{4}\sum_{\alpha}|{\boldsymbol{\phi}}^{\alpha}|^{4}.$ (9)
@@ II.4 $\Gamma$ point free energy
Under finite magnetic fields, the uniform quadrupole moment ${\boldsymbol{\phi}}_{0}^{\alpha}\equiv{\boldsymbol{\phi}}_{{\bm{k}}=\bf 0}^{\alpha}$ is induced via the effective field $\tilde{\bm{h}}$ in the second-order perturbation of the magnetic field $\bm{H}$. The free energy up to the fourth-order terms for the $\Gamma$ point modes are
$F_{\Gamma}$ $=\sum_{\alpha}\Big{(}-\tilde{\bm{h}}\cdot{\boldsymbol{\phi}}_{0}^{\alpha}+\frac{a_{\Gamma 0}}{2}|{\boldsymbol{\phi}}_{0}^{\alpha}|^{2}\Big{)}+J_{\Gamma 0}{\boldsymbol{\phi}}_{0}^{A}\cdot{\boldsymbol{\phi}}_{0}^{B}$
@@
$a_{\Gamma 0}$ $=T+12J_{2}+6J_{4},\quad J_{\Gamma 0}=4J_{1}+12J_{3}.$ (11)
The effective field is given by
@@
and $g=(g_{J}\mu_{\rm B})^{2}[7/(3E_{4})-1/E_{5}]$ with the excited $\Gamma_{4,5}$ CEF energy $E_{4,5}$.Hattori and Tsunetsugu (2014) $\mu_{\rm B}$ is the Bohr magneton and $g_{J}$ is the Lande’s $g$ factor for the multiplet with the angular momentum $J$; $g_{J}=4/5$ for the PrIr2Zn20 with $J=4$.
@@ II.5 Mode coupling terms
We now discuss mode coupling terms consisting of the fields at the L points and those at the $\Gamma$ points. The lowest-order contributions come from the cubic terms as
@@
$=-3\lambda_{b}\sum_{\alpha,n}\phi_{0}^{\alpha}\big{(}\phi_{n}^{\alpha}\big{)}^{2}\cos(\theta^{\alpha}_{0}+2\theta^{\alpha}_{n}).$ (13)
Here, we have parameterized
@@
and the coupling constant $\lambda_{b}$ has been introduced as an independent parameter. If $\lambda_{b}$ is estimated by Eq. (5), $\lambda_{b}=b$. The next leading order terms are in the fourth order and given by
@@
Again, the coupling constant $\lambda_{c}^{(1)}$ and $\lambda_{c}^{(2)}$ have been introduced as independent parameters. If $\lambda_{c}^{(1,2)}$ is estimated by Eq. (5), $\lambda^{(1)}_{c}=\lambda_{c}^{(2)}=c$.
In principle, ${\boldsymbol{\phi}}_{0}^{\alpha}$ can couple with ${\boldsymbol{\phi}}_{{\bm{k}}}^{\alpha}$ in the form of $\sim{\boldsymbol{\phi}}_{0}^{\alpha}{\boldsymbol{\phi}}_{\bm{k}}^{\alpha}{\boldsymbol{\phi}}_{-{\bm{k}}}^{\alpha}$ and ${\boldsymbol{\phi}}_{0}^{\alpha}{\boldsymbol{\phi}}_{0}^{\alpha}{\boldsymbol{\phi}}_{\bm{k}}^{\alpha}{\boldsymbol{\phi}}_{-{\bm{k}}}^{\alpha}$. In this paper, we retain the order parameter ${\boldsymbol{\phi}}_{n}^{\alpha}$ and the induced moment ${\boldsymbol{\phi}}_{0}^{\alpha}$ under magnetic fields since the other fields are not important for the phase transition with the L point fields. Some renormalization effects arising from fields at the X points will be discussed in Sec. IV and Appendix B.
@@ III Results: single-q orders
In this section, we carry out minimization of the free energy $F_{L}+F_{\Gamma}+\delta F_{3}+\delta F_{4}$ [Eqs. (6), (10), (13), and (15)], assuming the single-$\bm{q}$ orders. The purpose here is to understand how mode coupling effects affect the order parameter configuration. The results in this section will become bases for understanding the full phase diagram shown in Sec. IV, where we will take into account multiple-$\bm{q}$ configurations. Since we are interested in the phase transition owing to the order parameters at the L point, we assume the coefficients of the quadratic terms in the free energy at the other wavenumber are always positive. We first discuss the candidates of the stable configurations by analyzing the third- and fourth-order terms in $F_{\rm L}$. Then, we will show the numerical results of full temperature-field phase diagram.